OpenStudy (anonymous):

determine the value o k such that the first polynomial is divisible by the second k²x⁴-3kx²-4 : x-1

OpenStudy (anonymous):

as x-1 is a factor of the expression, place 1 in place of x in the polynomial and make it zero to find the value of k.. k^2-3k-4=0 (k+1)(k-4)=0 so, k=4,-1

OpenStudy (anonymous):

@TomasJoao , look at this ^^